User Tools

Site Tools


uc:uc20250701_july_1_2025

APO 3.5-m Users Committee Meeting, 07/01/2025


User feedback and comments from institutional representatives

  • BYU - nothing to report
  • Colorado - no report
  • Georgia State - nothing to report
  • JHU - nothing to report
  • NAPG - no report
  • NMSU - nothing to report
  • Oklahoma - nothing to report
  • Seattle - nothing to report
  • UVa - nothing to report
  • Washington - nothing to report
  • Wyoming - nothing to report in terms of 3.5m telescope usage, but informationally: Wyoming will be advertising a tenure-track astronomer position in the coming days. It will have an August 15 deadline for a January 2026 start date.

Telescope and Instruments Report

The detailed site report is included below.

3.5-m Telescope and Instruments Highlights, 6/03/25 – 07/01/25

1) Overview

The first half of June was rather dry. A large lightning strike caused a fire in difficult terrain, north of Cloudcroft on Mescalero tribal land, which reached over 800 acres. At the time of this report it is currently 97% contained. The last two weeks of June have seen a welcome relief in the dry conditions with the start of summer monsoonal weather patterns. The month of June had one visiting PI instrument scheduled but the team cancelled their travel and supported staff remotely. An REU group from U. Wyoming were additional visitors this month. Two large public outreach groups of high school and college summer science programs visited and were given tours. We had an inadvertent power cycle on two important UPS systems that power cycled several telescope subsystems including a few instruments. We were able to quickly recover.

As a reminder, summer shutdown is scheduled for July 7 - 17. Relatively short with only standard instrument and telescope maintenance.

2) Operations

3.5m Telescope: The telescope is working as expected. Seasonal motion errors have been infrequent. The secondary mirror was homed and the telescope positions reinitialized as part of the UPS power recovery mentioned above. A refinement in pointing was needed over the weekend as well.

0.5m Telescope: The telescope is working as expected. ACP Library support errors are still occurring even with the dcam-spare camera swap but with much lower frequency. The off-axis guider has not been very reliable, requiring power cycles.

KOSMOS: System is cooled and stable. A report of increased dark current for long exposures was received by a user. We hope to address this over summer shutdown by doing a vacuum servicing of the instrument.

ARCTIC: The diffuser rotation mechanism is still unreliable even after a full servicing. Troubleshooting is continuing. The mechanism that moves the diffuser in and out of the optical path is still functional in the meantime. The rest of the instrument is cooled. The CCD had an instance of a dropped quadrant as a result of the UPS power cycle.

Agile: The camera is non-operational; the thermoelectric cooler failed again. The camera is warm and we do not plan further repair work with this camera. We are planning to decommission the camera. The Agile instrument rotator is still not performing nominally and we are troubleshooting it further to prepare for SoonerCam.

ARCES: System is currently up and usable. The instrument seems to be consuming LN2 at a greater rate. We will likely try to address this over summer shutdown. Work continues on a replacement and modern ICC.

DIS: System is cooled but in an unknown state for science. Decommissioning plans have begun.

NICFPS: System is cooled and usable. The ICS software has needed multiple restarts over the past month. It is unclear at this time the cause.

TripleSpec: System is cooled and usable.

APOLLO: The system is continues experiencing some technical issues with the laser and thermal control at the time of this report. A team from NASA is on site this week to address the issues. Until repaired the instrument is not usable for laser ranging.


2025 Q3 3.5m scheduling

There are several open B-halves available in late July and early August, including dark time! Users should look at the schedule and follow the usual channels for requesting this time (i.e. email Russet, Amanda, Ben, Nancy and their institutional scheduler when submitting requests, and provide a proposal cover page if you don't already have a program scheduled for the current quarter).


2025 Q3 0.5m scheduling

There are two weeks open on the Q3 ARCSAT schedule in early August and early September. Users should look at the schedule and follow the usual channels for requesting this time (i.e. email Russet, Amanda, Ben, Nancy and their institutional scheduler when submitting requests, and provide a proposal cover page if you don't already have a program scheduled for the current quarter).


Status of new instrument initiatives

The Ocotillo purchase orders are largely in or are in the process of being generated and submitted. Procurement times are uncertain. The SoonerCam survey results will be collated soon.


Annual Request to Update Publication Record

Please send Nancy any references for 3.5m-related publications (anything published from Jan 1 2024 to now). The nominal deadline was June 30 but we have received relatively few responses so users are encouraged to complete this action item. See last month’s minutes for explanation of why these metrics are important. The listing should include refereed publications and student theses and dissertations (not conference abstracts unless they are peer reviewed) that made use of data acquired with the 3.5m and/or ARCSAT, published from 2024 to the present. If a paper has been submitted to a journal but has not yet completely gone through the review process, please include that as well, with the name of the journal and “submitted.”


Open Positions at APO

We have two vacancies on the 3.5m side that we are eager to fill!

1) Observing specialist - there are two postings for this; applicants should choose the one that best matches their education and experience level:

2) 3.5m Telescope Engineer - there will be two postings for this; applicants should choose the one that best matches their education and experience level:

These will be advertised in the July AAS Job Register but please share this announcement with anyone whom you think would be well qualified for these positions.


ACTION ITEMS

Open action items from previous meetings:

  • Nancy: get Mt. Cuba Foundation proposal for new KOSMOS grisms submitted. STATUS: CLOSED. Nancy submitted the proposal on 6/23/25.
  • UC reps: ask your users to complete the survey regarding needs/desires for a high speed imaging capability. The link to this survey was sent out to 35m-general on May 6, 2025 by Mukremin Kilic and it is also here: https://docs.google.com/forms/d/e/1FAIpQLSeJG6zQg7YtGxGhAtaKULHyzZVZe29uHCiegcHdM0edglN1gw/viewform Bill is planning to collate the results, but if there are too few responses now, the survey will stay open longer. STATUS: OPEN.
  • UC reps: ask your users to send Nancy the bibliographic information for any papers published in 2024-2025 that are resulting from 3.5m or ARCSAT data. STATUS: OPEN

New action items from this meeting:

  • UC reps: remind your users to review the Q3 schedule and request OPEN/DD time if they can make a good case for it. STATUS: OPEN.

All Other Business

None.


Next meeting

The next meeting will be on August 5, 2025 at 10:30 MDT.


uc/uc20250701_july_1_2025.txt · Last modified: 2025/07/22 02:31 by nchanove